If you invest $5,550 today at an interest rate of 6.75 percent, compounded daily, how much money will you have in your savings account in 7 years?

Round the answer to two decimal places.
Answer #1

The formula for compound interest:
A = P (1 + r/n) (nt)

A = the future value of the investment/loan, including interest
P = the principal investment amount (the initial deposit or loan amount)
r = the annual interest rate (decimal)
n = the number of times that interest is compounded per year
t = the number of years the money is invested or borrowed for

P = $5550. r = 6.75/100 = 0.07 (decimal). n = 365. t = 7.

If we plug these figures into the formula, we get the following:

A = 5550 (1 + 6.75 / 365) ^ (365(7) =

So, the investment balance after 7 years is $$8,901.81.